OpenStudy (anonymous):

(5t^2+t over t-1)^2 over means fraction... find the power @-------College Al…

jhonyy9 (jhonyy9):

(t(5t+1)/(t-1))2 =(t2(25t2+10t +1))/(t2-2t+1)

OpenStudy (anonymous):

5t^2+t over t-1)^2 ( t-1)^2=t^2-2t+1 now divide (5t^2+t +0) from t^2-2t+1 (5t^2+t +0)/(t^2-2t+1)=5+(11t-5)/(t-1)^2